# A note on internality of certain differential systems

### v1

@@ Abstract.
We prove two results, generalizing certain theorems by Jin and Moosa [1], on the internality of the system of differential equations
$x^{\prime}$ $=f(x)$
$y^{\prime}$ $=g(x)y,$
@@
Let $K$ be a differential field of characteristic zero. We fix a universal extension $U$ of $K$ and assume that all differential field extensions of $K$ considered in this article are contained in $U.$ For any differential field extension $M$ of $K,$ the subfield of constants of $M$ is denoted by $C_{M}.$ A generic solution of a system of first order differential equations
(A) $x^{\prime}$ $=f(x,y)$
$y^{\prime}$ $=g(x,y),$
where $f,g$ are rational functions in two variables over $K,$ is a tuple $(x,y)\in U^{2}$ such that $x^{\prime}=f(x,y),\ y^{\prime}=g(x,y)$ and that the field transcendence degree $\mathrm{tr.deg}(K(x,y)|K)=2.$
The system (A) is said to be almost internal to constants (respectively, internal to constants) if there there are positive integers $l$ and $m,$ constants $c_{1},\dots,c_{l}\in C_{U}$ and $m$ generic solutions $(x_{i},y_{i}),1\leq i\leq m,$ such that for every generic solution $(x,y)$ of (A), $x,y$ are algebraic over the field $K(S)$ (respectively $x,y$ belongs to the field $K(S)$) where $S=\{x_{i},y_{i},c_{p}\ |\ 1\leq i\leq m,1\leq p\leq l\}.$
Understanding internality of a system of equations offers insight into the structure of its solution fields. These concepts are known to be closely related to Kolchin’s strongly normal extensions and it has its orgin in a work of Rosenlicht [2, Proposition]. In [1], Jin and Moosa prove two theorems, which appears as Theorem A and Theorem B in their paper, on almost internality of the system
$x^{\prime}$ $=f(x)$
$y^{\prime}$ $=xy$
where $f$ is a rational function in one variable. In Theorem 3 and Theorem 4 of this article, we generalize and extend both the results of Jin and Moosa.
@@
Let $(x,y)$ be a generic solution of the system (A). If the system (A) is almost internal to constants, then there exists a differential field extension $M$ of $K$ such that $M$ and $K(x,y)$ are free over $K$ and $M(x,y)$ contains two $M$-algebraically independent constants. Furthermore, if NN is a differential field intermediate to $M(x,y)$ and $M$ with $\mathrm{tr.deg}(N|M)=1,$ then $\mathrm{tr.deg}(C_{N}|C_{M})=1.$
@@
Let $(x_{1},y_{1}),\dots,(x_{m},y_{m})$ be generic solutions of (A) and $c_{1},\dots,c_{l}$ be constants such that $x$ and $y$ are algebraic over $M(c_{1},\dots,c_{l}),$ where $M:=K(x_{1},y_{1},\dots,x_{m},y_{m}).$ Note that $M$ and $K(x,y)$ are free over $K.$ We show that $M(x,y)$ contains two $M$-algebraically independent constants.
To see this, let $c_{1},\dots,c_{r}$ be an $M$-transcendence base of constants for $M(c_{1},\dots,c_{l}).$ Since $\mathrm{tr.deg}(M(c_{1},\dots,c_{l},x,y)|M=r,$ we must have $\mathrm{tr.deg}(M(c_{1},\dots,c_{l},x,y)|M(x,y))=r-2$ and therefore, there is a nonzero polynomial $P\in M(x,y)[X_{1},\dots,X_{r}]$ such that $P(c_{1},\dots,c_{r})=0.$ Since $c_{1},\dots,c_{r}$ are constants, it follows that such a polynomial $P$ must belong to $C_{M(x,y)}[X_{1},\dots,X_{r}]$ ([3, Section 14, Theorem 2]). Since $\mathrm{tr.deg}(M(c_{1},\dots,c_{l})|M)=r,$ it also follows that one of the coefficients of $P$ must be a constant nonalgebraic over $M.$ If we call this constant $u_{1}$ and replace $M$ by $M(u_{1}),$ then a similar argument shows the existence of another constant $u_{2}\in M(x,y)$ which is nonalgebraic over $M(u_{1}).$ Observe that either $u_{1}$ or $u_{2}$ must belong to $M(x,y)\setminus M(x).$
Let $N$ be a differential field intermediate to $M(x,y)$ and $M$. Since $\mathrm{tr.deg}(N|M)=1,$ there is nonzero polynomial $P\in C_{N}[X_{1},X_{2}]$ such that $P(u_{1},u_{2})=0.$ Again, by the same argument as above, there is a constant $v\in N$ which is not algebraic over $M.$
@@
Let $f$ be a rational function in one variable over $K.$ The equation $x^{\prime}=f(x)$ is internal to constants if and only if $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1},a_{2}\in K.$
@@
Let $x,x_{1},\dots,x_{m}$ be generic solutions of $x^{\prime}=f(x)$ and suppose that $x\in K(x_{1},\dots,x_{m},c_{1},\dots,c_{l}),$ where $c_{1},\dots,c_{l}$ are constants. Consider the differential field $M=K(x_{1},\dots,x_{m}).$ From [4, Chapter 2, Corollary 2], we know that $M(x)$ is generated over $M$ as a field by a set of constants. Therefore, we may assume that $M(x)=M(c_{1},\dots,c_{l}).$ We first claim that there exists an algebraic extension $\tilde{M}$ of $M$ such that $\tilde{M}(x)=\tilde{M}(c)$ for some constant $c$.
Note that $C_{M(x)}$ and $M$ are linearly disjoint over $C_{M}$ ([4, Chapter 2, Corollary 1]). Therefore, $C_{M}(c_{1},\dots,c_{l})$ and $M$ are linearly disjoint over $C_{M}.$ Since $M(c_{1},\dots,c_{l})=M(x),$ we have $\mathrm{tr.deg}(M(c_{1},\dots,c_{l})|M)=1$ and from linear disjointness, we obtain that $\mathrm{tr.deg}(C_{M}(c_{1},\dots,c_{l})|C_{M})=1.$ Now since $M(c_{1},\dots,c_{l})=M(x)$ is a rational function field, there is a finite algebraic extension $\tilde{C}_{M}$ of $C_{M}$ such that $\tilde{C}_{M}(c_{1},\dots,c_{l})=\tilde{C}_{M}(c)$ ([5, Theorem 5]). Any extension of a zero derivation on a field to its algebraic closure is again a zero derivation. Therefore, $c$ must also be a constant. Let $\tilde{M}:=M\tilde{C}_{M}$ and observe that
@@
Since $\tilde{M}(x)=\tilde{M}(c),$ there exists $\alpha,\beta,\gamma,\delta\in\tilde{M}$ such that
@@
To prove the converse, let $x,x_{1},x_{2},x_{3}$ be generic solutions of $x^{\prime}=a_{2}x^{2}+a_{1}x+a_{0}$ over $K.$ It is observed in [6, Page 102] that
@@
is a constant [^1] It is easily seen that the first order homogeneous equation $y^{\prime}=(a_{1}(x+x_{1}+x_{2}+x_{3})+a_{0})y$ has $(x-x_{2})(x_{3}-x_{1})$ and $(x-x_{1})(x_{3}-x_{2})$ as its solutions and therefore $c,$ being the ratio of these two solutions, must be a constant.. Since
@@
we now obtain $x\in K(x_{1},x_{2},x_{3},c).$ This proves that the generic type $x^{\prime}=a_{2}x^{2}+a_{1}x+a_{0}$ over $K$ is internal to constants. ∎
@@
(cf. [1, Theorem A]) Let $K$ be a differential field and $f$ be a nonzero rational function in one variable over $K$. The system
(B) $x^{\prime}$ $=f(x)$
$y^{\prime}$ $=xy$
is almost internal to constants and $x^{\prime}=f(x)$ is internal to constants if and only if $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1}\in K$ and $a_{2}$ is a nonzero rational number.
@@
We know from Theorem 2 that $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1},a_{2}\in K.$ Suppose that the system (B) is almost internal to constants having a generic solution $(x,y)$. Then by Proposition 1, there exists a differential field extension $M$ of $K$ such that $M(x,y)$ contains two $M$-algebraically independent constants. Then $M(x,y)$ contains a constant which is not in $M(x)$ and therefore the latter must contain an element $z$ such that $z^{\prime}=nxz$ for some positive integer $n$ ([7, Example 1.11]). Let $z=P/Q$ where $P,Q\in M[x].$ We can assume that $P$ and $Q$ are relatively prime and $Q$ is monic. Write $P=\alpha_{p}x^{p}+\alpha_{p-1}x^{p-1}+\cdots,Q=x^{q}+b_{q-1}x^{q-1}+\cdots\in M[x]$ with $\alpha_{p}\neq 0.$ Then
@@
and we obtain that $(p-q)a_{2}=n.$ Since $n\neq 0,$ $a_{2}=n/(p-q)$ is a nonzero rational number.
For the converse, suppose that $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1}\in K$ and $a_{2}=m_{1}/m_{2},$ $(m_{1},m_{2})\in\mathbb{Z}^{2}$ a nonzero rational number. Let $(x,y)$ be a generic solution of the system (B). We shall now find generic solutions $(x_{1},y_{1})$ and $(x_{2},y_{2})$ of the system and constants $c_{1},c_{2}$ so that $x,y$ are algebraic over the field $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ This will then complete the proof of the theorem.
Let $v$ be an element algebraic over the field $K(x,y)$ satisfying $v^{m_{2}}=y^{-m_{1}}.$ Then $v^{\prime}=-a_{2}xv,$ $(a_{2}x)^{\prime}=a_{2}x^{\prime}=(a_{2}x)^{2}+a_{1}(a_{2}x)+a_{2}a_{0}$ and $v^{\prime\prime}=a_{1}v^{\prime}-a_{2}a_{1}v.$ Consider the rational function field $K(v,v^{\prime})(v_{1},v^{\prime}_{1},v_{2},v^{\prime}_{2})$ and extend the derivation on $K(v,v^{\prime})$ to $K(v,v^{\prime})(v_{1},v_{2},v^{\prime}_{1},v^{\prime}_{2})$ by declaring $v^{\prime\prime}_{1}=a_{1}v^{\prime}_{1}-a_{2}a_{1}v_{1}$ and $v^{\prime\prime}_{2}=a_{1}v^{\prime}_{2}-a_{2}a_{1}v_{2}$. Let $N=K(v_{1},v^{\prime}_{1},v_{2},v^{\prime}_{2}).$ Note that $v_{1},v_{2}$ are $C_{U}$-linearly independent and that the Wronskian of $v,v_{1},v_{2}$ is zero. Therefore $v=c_{1}v_{1}+c_{2}v_{2},$ for some constants $c_{1},c_{2}.$ This implies, $v\in N(c_{1},c_{2})$ and since $a_{2}x=-v^{\prime}/v\in N(c_{1},c_{2}),$ we obtain that $x\in N(c_{1},c_{2}).$
Let $y_{1}$ and $y_{2}$ be elements algebraic over the field $N$ such that $v^{m_{2}}_{1}=y^{-m_{1}}_{1}$ and $v^{m_{2}}_{2}=y^{-m_{1}}_{2}.$ Let $x_{1}:=-v^{\prime}_{1}/a_{2}v_{1}$ and $x_{2}:=-v^{\prime}_{2}/a_{2}v_{2}.$ Then, it is easy to see that $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are generic solutions of the system (B). Now we consider the field $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ Clearly, $v_{1}$ and $v_{2}$ are algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2})$ and since $v=c_{1}v_{1}+c_{2}v_{2},$ we obtain that $v$ is also algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ Since $x=-v^{\prime}/a_{2}v,$ we obtain that $x$ is algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ Finally, $v^{m_{2}}=y^{-m_{1}}$ implies that $y$ is algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$
@@
(cf. [1, Theorem B]) Let $C$ be an algebraically closed differential field with the zero derivation and $f,g$ be two nonzero rational functions in one variable over $C.$ The system
(C) $x^{\prime}$ $=f(x)$
$y^{\prime}$ $=g(x)y$
is almost internal to constants if and only if the following two conditions hold:
@@
    $\frac{1}{f(x)}=\frac{\partial u}{\partial x}$ or $\frac{1}{f(x)}=\frac{c}{u}\frac{\partial u}{\partial x}$ for some $c\in C\setminus\{0\}$ and $u\in C(x).$
@@
Let $(x,y)$ be a generic solution of the system (C) and $F=C(x,y).$ We first make the following observations. Let $z\in F\setminus C$ such that $z^{\prime}=0.$ Then $z\notin C(x);$ otherwise $f(x)\frac{\partial z}{\partial x}=0,$ which is a contradiction. Therefore, $z\in F\setminus C(x).$ As noted in the proof of Theorem 3, there exists a positive integer $m$ and an element $v\in C(x)$ such that $v^{\prime}=mg(x)v.$ Therefore, $f(x)\frac{\partial v}{\partial x}=mg(x)v$ and we thus obtain
@@
Along with these observations we suppose that the system (C) is almost internal to constants. Then, by Proposition 1, there exists a differential field extension $M$ of $C$ such that $M(x,y)$ contains two $M$-algebraically independent constants. Furthermore, since $M(x)$ is a differential subfield of $M(x,y)$ with $\mathrm{tr.deg}(M(x)|M)=1,$ it follows that $\mathrm{tr.deg}(C_{M(x)}|C_{M})=1.$ By [2, Proposition], $C(x)$ contains an element $u$ such that either $u^{\prime}=1$ or $u^{\prime}=cu$ for some nonzero $c\in C.$ Then $\frac{1}{f(x)}$ is either $\frac{\partial u}{\partial x}$ or $\frac{1}{f(x)}=\frac{c}{u}\frac{\partial u}{\partial x}.$ This implies that condition (i) holds.
To prove that condition (ii) also holds, we consider the following two cases. Suppose that $C_{F}\neq C$. Then by the observation made in the first paragraph, the condition (ii) holds. Now suppose that $C_{F}=C.$ Since there exist a differential field $M$ such that $M$ and $F$ are free over $C$ and $MF$ contains two $M$-algebraically independent constants, by [8, Proposition 4.2], there is a differential field $L$ intermediate to $C$ and $F$ such that $\mathrm{tr.deg}(L|C)=2$ and $L$ can be embedded in a strongly normal extension of $C.$ Since $L$ is contained in the purely transcendental extension $C(x,y),$ by [8, Theorem 1.2], $L$ can be embedded in a Picard-Vessiot extension of $C.$ Then $L=C(u,v)$ where $u,v$ are $C$-algebraically independent, $v^{\prime}=c_{2}v$ for some nonzero $c_{2}\in C$ and either $u^{\prime}=1$ or $u^{\prime}=c_{1}u$ for some nonzero $c_{1}\in C.$ As $C(x)$ already has an element $z$ such that $z^{\prime}=1$ or $z^{\prime}=cz,$ it is guaranteed that there exists an element $\zeta\in C(x,y)$ (either $u$ or $v$) such that $\zeta^{\prime}=a\zeta$ for some nonzero $a\in C.$ By the Kolchin-Ostrowski theorem, there exists integers $m,n$ such that $h:=y^{m}\zeta^{n}\in C(x).$ Then
@@
From the above equation we obtain

### v2

@@ Abstract.
We prove two results, generalizing certain theorems by Jin and Moosa [JM20], on the internality of the system of differential equations
[…]
@@
Let $K$ be a differential field of characteristic zero. We fix a universal differential field extension [^1] A differential field extension $U$ of $K$ is said to be an universal extension of $K$ if for any finitely generated differential field extension of $F$ of $K$ contained in $U$ and for any finitely generated differential field extension $F_{1}$ of $F,$ which is not necessarily contained in $U$, there exists a differential $F$-embeddeding from $F_{1}$ into $U$ (see [Kol73, page 133]). $U$ of $K$ and assume that all differential field extensions of $K$ considered in this article are contained in $U.$ For any differential field extension $M$ of $K,$ the subfield of constants of $M$ is denoted by $C_{M}.$ A generic solution over $K$ of a system of first order differential equations
[…]
where $f_{1},\dots,f_{n}$ are rational functions in $n$ variables over $K,$ is a tuple $(y_{1},\dots,y_{n})\in U^{n}$ such that for each $i,1\leq i\leq n,$ $y^{\prime}_{i}=f_{i}(y_{1},\dots,y_{n})$ and the field transcendence degree $\mathrm{tr.deg}(K(y_{1},\dots,y_{n})|K)=n.$
[…]
The model theoretic concept of internality of a system of differential equations is known to be closely related to Kolchin’s strongly normal extensions (see [Poi01, Section 2.7]), and it has its origin in the works of Rosenlicht [Ros74, Proposition] and Zilber [Zil77]. In [JM20], Jin and Moosa prove two theorems, which appear as Theorem A and Theorem B in their paper, on almost internality of the system of differential equations
[…]
where $f$ is a rational function in one variable. In Theorems 4 and 5 of this article, we generalize and extend both these results of Jin and Moosa. In particular, Theorem 4 completely answers the question of when the system (L) is almost internal to the constants, given that $x^{\prime}=f(x)$ is internal to the constants.
@@
Let
[…]
be a system almost internal to the constants. Then, there exists a finitely generated differential field extension $M$ of $K$ such that for any generic solution $(x,y)$ over $M$ of the system [^2] Generic solutions over $M$ of the system always exist. To see this, consider the rational function field $M(s,t)$ and equip this field with the derivation induced by the system: $s^{\prime}=f(s,t)$ and $t^{\prime}=g(s,t)$. Then, by definition, there is a differential $M$-embedding of $M(s,t)$ into $U.$ The image of $s,t$ under this embedding provides a generic solution over $M$ of the system., the differential field $M(x,y)$ contains two $M$-algebraically independent constants. Furthermore, if $N$ is a differential field intermediate to $M(x,y)$ and $M$ with $\mathrm{tr.deg}(N|M)=1,$ then $\mathrm{tr.deg}(C_{N}|C_{M})=1.$
@@
Since the system is almost internal to the constants, there are generic solutions $(x_{1},y_{1}),\dots,(x_{m},y_{m})$ over $K$ of the system such that for any generic solution $(x,y)$ of the system over $M:=K(x_{1},y_{1},\dots,x_{m},y_{m}),$ there are constants $c_{1},\dots,c_{l}$ such that $x$ and $y$ are algebraic over $M(c_{1},\dots,c_{l}).$ We now show that $M(x,y)$ contains two $M$-algebraically independent constants. Let $c_{1},\dots,c_{r}$ be an $M$-transcendence basis of constants for $M(c_{1},\dots,c_{l}).$ Since $\mathrm{tr.deg}(M(c_{1},\dots,c_{l},x,y)|M)=r,$ we must have $\mathrm{tr.deg}(M(c_{1},\dots,c_{l},x,y)|M(x,y))=r-2.$ Therefore, there is a nonzero polynomial $P\in M(x,y)[X_{1},\dots,X_{r}]$ such that $P(c_{1},\dots,c_{r})=0.$ Since $c_{1},\dots,c_{r}$ are constants, it follows that such a polynomial $P$ must belong to $C_{M(x,y)}[X_{1},\dots,X_{r}]$ ([Kol48, Section 14, Theorem 2]). As $\mathrm{tr.deg}(M(c_{1},\dots,c_{l})|M)=r,$ it also follows that one of the coefficients of $P$ must be a constant nonalgebraic over $M.$ If we call this constant $u_{1}$ and replace $M$ by $M(u_{1}),$ then a similar argument shows the existence of another constant $u_{2}\in M(x,y),$ which is nonalgebraic over $M(u_{1}).$ Observe that either $u_{1}$ or $u_{2}$ must belong to $M(x,y)\setminus M(x).$
Let $N$ be a differential field intermediate to $M(x,y)$ and $M$. Since $\mathrm{tr.deg}(N|M)=1,$ there is a nonzero polynomial $P\in C_{N}[X_{1},X_{2}]$ such that $P(u_{1},u_{2})=0.$ Again, by the same argument as above, there is a constant $v\in N$ which is not algebraic over $M.$
@@
Let $f$ be a rational function in one variable over $K.$ The equation $x^{\prime}=f(x)$ is internal to the constants if and only if $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1},a_{2}\in K.$
@@
Let $x^{\prime}=f(x)$ be internal to the constants. Then, there are generic solutions $x_{1},\dots,x_{m}$ over $K$ such that for any generic solution $x$ over $M:=K(x_{1},\dots,x_{m})$ there are constants $c_{1},\dots,c_{l}$ such that $x\in M(c_{1},\dots,c_{l})$. From [Kol73, Chapter 2, Corollary 2], we know that $M(x)$ is generated over $M$ as a field by a set of constants. Therefore, we may assume that $M(x)=M(c_{1},\dots,c_{l}).$ We first claim that there exists an algebraic extension $\tilde{M}$ of $M$ such that $\tilde{M}(x)=\tilde{M}(c)$ for some constant $c$.
Note that $C_{M(x)}$ and $M$ are linearly disjoint over $C_{M}$ ([Kol73, Chapter 2, Corollary 1]). Therefore, $C_{M}(c_{1},\dots,c_{l})$ and $M$ are linearly disjoint over $C_{M}.$ Since $M(c_{1},\dots,c_{l})=M(x),$ we have $\mathrm{tr.deg}(M(c_{1},\dots,c_{l})|M)=1$ and from linear disjointness, we obtain $\mathrm{tr.deg}(C_{M}(c_{1},\dots,c_{l})|C_{M})=1.$ Now since $M(c_{1},\dots,c_{l})=M(x)$ is a rational function field, there is a finite algebraic extension $\tilde{C}_{M}$ of $C_{M}$ such that $\tilde{C}_{M}(c_{1},\dots,c_{l})=\tilde{C}_{M}(c)$ ([Che63, Theorem 5]). Any extension of a zero derivation on a field to its algebraic closure is again a zero derivation. Therefore, $c$ must also be a constant. Let $\tilde{M}:=M\tilde{C}_{M}$ and observe that
@@
Since $\tilde{M}(x)=\tilde{M}(c),$ there exists $\alpha,\beta,\gamma,\delta\in\tilde{M}$ with $\alpha\delta-\beta\gamma\neq 0$ such that
@@
To prove the converse, let $x,x_{1},x_{2},x_{3}$ be distinct generic solutions of $x^{\prime}=a_{2}x^{2}+a_{1}x+a_{0}$ over $K.$ It is observed in [Nag20, Page 102] that
@@
is a constant [^3] It is easily seen that the first order differential equation $y^{\prime}=(a_{1}(x+x_{1}+x_{2}+x_{3})+a_{0})y$ has $(x-x_{2})(x_{3}-x_{1})$ and $(x-x_{1})(x_{3}-x_{2})$ as its solutions and therefore $c,$ being the ratio of these two solutions, must be a constant.. Since
@@
we now obtain $x\in K(x_{1},x_{2},x_{3},c).$ This proves that the equation $x^{\prime}=a_{2}x^{2}+a_{1}x+a_{0}$ is internal to the constants. ∎
Remark 3.
We note here that a particular case of Theorem 2 can be derived from certain recent results ([JM25, KRS24]). Suppose that $K$ is an algebraically closed field and that the equation $x^{\prime}=f(x)$ is both internal to the constants and weakly orthogonal to the constants [^4] That is, the rational function field $K(x)$ with the extended derivation $x^{\prime}=f(x)$ has the same field of constants as $K.$. Then, it is known that the binding group of the equation $x^{\prime}=f(x)$ is a linear algebraic group. Therefore, one can apply [JM25, Proposition 6.5] and conclude that $K(x)=K(y),$ where $y^{\prime}=ay^{2}+by+c$ for some $a,b,c\in K.$ Writing
$x=\frac{\alpha y+\beta}{\gamma y+\delta},$
where $\alpha,\beta,\gamma,\delta\in K$ with $\alpha\delta-\beta\gamma\neq 0,$ and taking derivatives, we obtain the conclusion of Theorem 2 that $f$ is a polynomial over $K$ of degree at most $2.$
The conditions that $x^{\prime}=f(x)$ is both internal to the constants and weakly orthogonal to the constants imply that the differential field $K(x)$ can be embedded in a strongly normal extension of $K.$ Now, applying [KRS24, Theorem 1.1 (iii)], we obtain that $f$ is a polynomial over $K$ of degree at most $2.$ Thus, Theorem 2 provides a generalization of these results, as it requires neither that $x^{\prime}=f(x)$ be weakly orthogonal to the constants nor that $K$ be an algebraically closed field.
Under the assumption that $x^{\prime}=f(x)$ is internal to the constants with the binding group not of dimension $3,$ it is shown in [JM20, Theorem A] that the system (L) is almost internal to the constants if and only if it splits; which implies, the system can be transformed into a system of the form
$\begin{cases}\begin{aligned} x^{\prime}&=f(x)\\ y^{\prime}&=ay\\ \end{aligned}\end{cases},$
where $a\in K.$ In the next theorem, using Theorem 2 and Proposition 1 and without any assumption on the binding group of $x^{\prime}=f(x)$, we shall completely determine when the system (L) is almost internal to the constants.
@@
Let $K$ be a differential field and $f$ be a nonzero rational function in one variable over $K$. The system (L) is almost internal to the constants and $x^{\prime}=f(x)$ is internal to the constants if and only if $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1}\in K$ and $a_{2}$ is a nonzero rational number.
@@
We know from Theorem 2 that $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1},a_{2}\in K.$ Suppose that the system (L) is almost internal to the constants. By Proposition 1, there exists a differential field extension $M$ of $K$ such that for any generic solution $(x,y)$ over $M,$ the differential field $M(x,y)$ contains two $M$-algebraically independent constants. Then $M(x,y)$ contains a constant which is not in $M(x)$ and therefore the latter must contain an element $z$ such that $z^{\prime}=nxz$ for some positive integer $n$ ([Mag94, Example 1.11]). Let $z=P/Q$ where $P,Q\in M[x].$ We can assume that $P$ and $Q$ are relatively prime and $Q$ is monic. Write $P=\alpha_{p}x^{p}+\alpha_{p-1}x^{p-1}+\cdots,Q=x^{q}+b_{q-1}x^{q-1}+\cdots\in M[x]$ with $\alpha_{p}\neq 0.$ Then
@@
and we obtain $(p-q)a_{2}=n.$ Since $n\neq 0,$ $a_{2}=n/(p-q)$ is a nonzero rational number.
For the converse, suppose that $f(x)=a_{2}x^{2}+a_{1}x+a_{0},$ where $a_{0},a_{1}\in K$ and $a_{2}=m_{1}/m_{2},$ $(m_{1},m_{2})\in\mathbb{Z}^{2}$ a nonzero rational number. Let $(x,y)$ be a generic solution of the system (L). We shall now find generic solutions $(x_{1},y_{1})$ and $(x_{2},y_{2})$ of the system and constants $c_{1},c_{2}$ so that $x,y$ are algebraic over the field $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ This will then complete the proof of the theorem.
Let $v$ be an element algebraic over the field $K(x,y)$ satisfying $v^{m_{2}}=y^{-m_{1}}.$ Then $v^{\prime}=-a_{2}xv,$ $(a_{2}x)^{\prime}=a_{2}x^{\prime}=(a_{2}x)^{2}+a_{1}(a_{2}x)+a_{2}a_{0}$ and $v^{\prime\prime}=a_{1}v^{\prime}-a_{2}a_{0}v.$ Consider the rational function field $K(v,v^{\prime})(v_{1},v^{\prime}_{1},v_{2},v^{\prime}_{2})$ and extend the derivation on $K(v,v^{\prime})$ to $K(v,v^{\prime})(v_{1},v_{2},v^{\prime}_{1},v^{\prime}_{2})$ by declaring $v^{\prime\prime}_{1}=a_{1}v^{\prime}_{1}-a_{2}a_{0}v_{1}$ and $v^{\prime\prime}_{2}=a_{1}v^{\prime}_{2}-a_{2}a_{0}v_{2}$. Let $N=K(v_{1},v^{\prime}_{1},v_{2},v^{\prime}_{2}).$ Note that $v_{1},v_{2}$ are $C_{U}$-linearly independent and that the Wronskian of $v,v_{1},v_{2}$ is zero. Therefore $v=c_{1}v_{1}+c_{2}v_{2},$ for some constants $c_{1},c_{2}.$ This implies, $v\in N(c_{1},c_{2})$ and since $a_{2}x=-v^{\prime}/v\in N(c_{1},c_{2}),$ we obtain that $x\in N(c_{1},c_{2}).$
Let $y_{1}$ and $y_{2}$ be elements algebraic over the field $N$ such that $v^{m_{2}}_{1}=y^{-m_{1}}_{1}$ and $v^{m_{2}}_{2}=y^{-m_{1}}_{2}.$ Let $x_{1}:=-v^{\prime}_{1}/a_{2}v_{1}$ and $x_{2}:=-v^{\prime}_{2}/a_{2}v_{2}.$ Then, it is easy to see that $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are generic solutions of the system (L). Now we consider the field $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ Clearly, $v_{1}$ and $v_{2}$ are algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2})$ and since $v=c_{1}v_{1}+c_{2}v_{2},$ we obtain that $v$ is also algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ Since $x=-v^{\prime}/a_{2}v,$ we obtain that $x$ is algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$ Finally, $v^{m_{2}}=y^{-m_{1}}$ implies that $y$ is algebraic over $K(y_{1},x_{1},y_{2},x_{2},c_{1},c_{2}).$
@@
Let $C$ be an algebraically closed differential field with zero derivation, and $f,g$ be two nonzero rational functions in one variable over $C.$ The system
[…]
is almost internal to the constants if and only if the following two conditions hold:
@@
    $\frac{1}{f(x)}=\frac{\partial u}{\partial x}$ or $\frac{1}{f(x)}=\frac{c}{u}\frac{\partial u}{\partial x}$ for some $c\in C\setminus 0$ and $u\in C(x).$
@@
Let $(x,y)$ be a generic solution of the system (A) and $F=C(x,y).$ We first make the following observations. Let $z\in F\setminus C$ such that $z^{\prime}=0.$ Then $z\notin C(x);$ otherwise $f(x)\frac{\partial z}{\partial x}=0,$ which is a contradiction. Therefore, $z\in F\setminus C(x).$ As noted in the proof of Theorem 4, there exists a positive integer $m$ and an element $v\in C(x)$ such that $v^{\prime}=mg(x)v.$ Therefore, $f(x)\frac{\partial v}{\partial x}=mg(x)v$ and we obtain
@@
Along with these observations, we suppose that the system (A) is almost internal to the constants. Then, by Proposition 1, there exists a differential field extension $M$ of $C$ such that $M(x,y)$ contains two $M$-algebraically independent constants. Furthermore, since $M(x)$ is a differential subfield of $M(x,y)$ with $\mathrm{tr.deg}(M(x)|M)=1,$ it follows that $\mathrm{tr.deg}(C_{M(x)}|C_{M})=1.$ By [Ros74, Proposition], $C(x)$ contains an element $u$ such that either $u^{\prime}=1$ or $u^{\prime}=cu$ for some nonzero $c\in C.$ Then $\frac{1}{f(x)}$ is either $\frac{\partial u}{\partial x}$ or $\frac{1}{f(x)}=\frac{c}{u}\frac{\partial u}{\partial x}.$ This implies that condition (i) holds.
To prove that condition (ii) also holds, we consider the following two cases. Suppose that $C_{F}\neq C$. Then, by the observation made in the first paragraph, the condition (ii) holds. Now suppose that $C_{F}=C.$ Since there exists a differential field $M$ such that $M$ and $F$ are free over $C$ and $MF$ contains two $M$-algebraically independent constants, by [KS25, Proposition 4.2], there is a differential field $L$ intermediate to $C$ and $F$ such that $\mathrm{tr.deg}(L|C)=2$ and $L$ can be embedded in a strongly normal extension of $C.$ Since $L$ is contained in the purely transcendental extension $C(x,y),$ by [KS25, Theorem 1.2], $L$ can be embedded in a Picard-Vessiot extension of CC [^5] The existence of the field $L$ can also be obtained using results from model theory (see [EJ24, Theorem 3.5]). The differential Galois group of any Picard-Vessiot extension of $C$ is a connected commutative linear algebraic group over $C$. Therefore, by the fundamental theorem of differential Galois theory, $L$ must also be a Picard-Vessiot extension of $C$ with a differential Galois group isomorphic to $\mathbb{G}_{a}\times\mathbb{G}_{m}$ or $\mathbb{G}_{m}\times\mathbb{G}_{m}.$ This implies, $L=C(\alpha,\beta)$ where $\alpha,\beta$ are $C$-algebraically independent, $\beta^{\prime}=c_{2}\beta$ for some nonzero $c_{2}\in C$ and either $\alpha^{\prime}=1$ or $\alpha^{\prime}=c_{1}\alpha$ for some nonzero $c_{1}\in C.$ Consequently, there exists an element $\zeta\in C(x,y)$ such that $\zeta^{\prime}=a\zeta$ for some nonzero $a\in C.$ By the Kolchin-Ostrowski theorem, there exists integers $m,n$ such that $h:=y^{m}\zeta^{n}\in C(x).$ Then
@@
From the above equation, we obtain